THICK SUBCATEGORIES OF MODULES OVER COMMUTATIVE

RINGS

(WITH AN APPENDIX BY SRIKANTH IYENGAR)

HENNING KRAUSE

Abstract. For a commutative noetherian ring A, we compare the support of a complex of A-modules with the support of its cohomology. This leads to a classification of all full subcategories of A-modules which are thick (that is, closed under taking kernels, cokernels, and extensions) and closed under taking direct sums.

1. Introduction

Let A be a commutative noetherian ring. We consider the category Mod A of Amodules and the spectrum Spec A of prime ideals of A. Given a complex X of A-modules, we wish to relate its support (in the sense of Foxby [6])

Supp X = {𝔭 ∈ Spec A | X ⊗_{A}^{L} k(𝔭) ≠ 0}

to the support of its cohomology

Supp H^{∗}X = ⋃_{i∈ℤ}Supp H^{i}X.

In some cases we have the equality

Supp X = Supp H^{∗}X,

for example when A is a Dedekind domain or when H^{∗}X is finitely generated over A. The failure of this equality is the main theme of recent joint work with Benson and Iyengar [2], which is motivated by the study of support varieties of modular representations. In this paper we establish a closely related classification of thick subcategories of Mod A and prove the following result.

Theorem 1.1. For a subset Φ of Spec A the following conditions are equivalent: (1) For every complex X of A-modules we have

Supp X ⊆ Φ ⟺ Supp H^{∗}X ⊆ Φ.

(2) The A-modules M with Supp M ⊆ Φ form a thick subcategory of Mod A.

(3) Every map I^{0} → I^{1} between injective A-modules with Ass I^{i} ⊆ Φ (i = 0, 1) can be completed to an exact sequence I^{0} → I^{1} → I^{2} such that I^{2} is injective and Ass I^{2} ⊆ Φ.

Recall that a classical result of Gabriel [8] provides a bijection between the set of localizing subcategories of Mod A and the set of specialization closed subsets of Spec A. More recently, a number of authors studied subcategories of Mod A in terms of subsets of Spec A; see [9, 11, 16]. In this paper we generalize Gabriel’s result in two directions.

Version from March 4, 2007.

The first direction (via associated primes) is fairly elementary but seems to be new. The second direction (via support) leads to a classification of thick subcategories of Mod A and corrects a result of Hovey in [11]. This classification is formulated in terms of subsets Φ ⊆ Spec A satisfying the equivalent conditions of Theorem 1.1. We call such subsets coherent and establish the following.

Theorem 1.2. For a commutative noetherian ring A the following conditions are equivalent:

(1) The Krull dimension of A is at most one.

(2) Every subset of Spec A is coherent.

(3) Supp X = Supp H^{∗}X for every complex X of A-modules.

The proof of this theorem is illustrated by some explict examples of subsets of Spec A which are not coherent. It would be interesting to have a geometric interpretation of coherent subsets in terms of the Zariski topology on Spec A.

2. Subcategories via associated primes

Let M be an A-module. Recall that a prime ideal 𝔭 is associated to M, if A/𝔭 is isomorphic to a submodule of M. We denote by Ass M the set of all prime ideals which are associated to M.

Theorem 2.1. The map sending a subcategory C of Mod A to

Ass C = ⋃_{M∈C} Ass M

induces a bijection between the set of full subcategories of Mod A, which are closed under taking submodules, extensions, and direct unions, and the set of subsets of Spec A. The inverse map sends a subset Φ of Spec A to

Ass^{−1} Φ = {M ∈ Mod A | Ass M ⊆ Φ}.

The proof uses some basic facts about associated primes and the structure of injective modules. The injective envelope of a module M is denoted by E(M) and we observe that Ass E(M) = Ass M.

Lemma 2.2. Let M be an A-module. Given a submodule N ⊆ M and a family (M_{i}) of submodules satisfying M = ⋃_{i} M_{i}, we have

Ass N ⊆ Ass M ⊆ Ass N ∪ Ass M/N and Ass M = ⋃_{i} Ass M_{i}.

Proof. See [4, Chap. IV, §1]. □

For a prime ideal 𝔭 we denote by k(𝔭) its residue field.

Lemma 2.3. Let I be an indecomposable injective A-module and 𝔭 its associated prime ideal. Then I is obtained from A/𝔭 by taking extensions and direct unions. More precisely,

(1) I is obtained from k(𝔭) by taking extensions and direct unions, and

(2) k(𝔭) is a direct union of copies of A/𝔭.

Proof. We sketch the argument and refer to [5, Chap. X, §8] for details. For each integer n ≥ 0 let I_{n} denote the submodule of I consisting of all elements annihilated by 𝔭^{n}. Then we have I = ⋃_{n≥0} I_{n} and each factor I_{n+1}/I_{n} is isomorphic to a finite direct sum of copies of the residue field k(𝔭). Now observe that k(𝔭) is the field of fractions of A/𝔭 and therefore a direct union of the form

k(𝔭) = ⋃_{0≠x∈A/𝔭}x^{−1}A/𝔭.  □

Proof of Theorem 2.1. Let Φ be a subset of Spec A. Then the subcategory Ass^{−1} Φ is closed under taking submodules, extensions, and direct unions, by Lemma 2.2. Clearly, we have

Ass(Ass^{−1} Φ) = Φ.

Now let C be a subcategory of Mod A, which is closed under taking submodules, extensions, and direct unions. We claim that

Ass^{−1}(Ass C) = C.

The inclusion Ass^{−1}(Ass C) ⊇ C is clear. Now suppose that M is a module contained in Ass^{−1}(Ass C). Then its injective envelope E(M) is a direct sum of copies of the form E(A/𝔭) with 𝔭 ∈ Ass C, since Ass E(M) = Ass M. But 𝔭 ∈ Ass C implies A/𝔭 ∈ C, and therefore E(A/𝔭) belongs to C, by Lemma 2.3. It follows that E(M) belongs to C and therefore M ∈ C. This finishes the proof. □

We state a number of consequences of Theorem 2.1.

Corollary 2.4. For a full subcategory C of Mod A the following conditions are equivalent.

(1) C is closed under taking submodules, extensions, and direct unions.

(2) There exists a subset Φ of Spec A such that C consists of all A-modules M satisfying Ass M ⊆ Φ.

(3) There exists an injective A-module I such that C consists of all A-modules which admit a monomorphism into a direct sum of copies of I.

Proof. (1) ⇔ (2): This is an immediate consequence of Theorem 2.1.

(2) ⇒ (3): Take I = ⨁_{𝔭∈Φ} E(A/𝔭). Then Ass M ⊆ Φ for every submodule M of a direct sum of copies of I. On the other hand, if Ass M ⊆ Φ, then Ass E(M) ⊆ Φ and therefore E(M) is a submodule of a direct sum of copies of I.

(3) ⇒ (1): Clear. □

Next we restrict the map C ↦ Ass C to the category mod A of all finitely generated A-modules.

Corollary 2.5 ([16, Theorem 4.1]). The map D ↦ Ass D induces a bijection between the set of full subcategories of mod A, which are closed under taking submodules and extensions, and the set of subsets of Spec A.

Proof. Consider the map C ↦ C ∩ mod A between

(i) the set of full subcategories of Mod A, which are closed under taking submodules, extensions, and direct unions, and

(ii) the set of full subcategories of mod A, which are closed under taking submodules and extensions.

This map is bijective; its inverse sends a subcategory D from (ii) to the full subcategory of Mod A consisting of all direct unions of modules in D. The composition of the first map C ↦ C∩mod A with D ↦ Ass D is the bjection from Theorem 2.1. Thus D ↦ Ass D is bijective. □

Recall that a full subcategory C of Mod A is localizing if C is closed under taking submodules, factor modules, extensions, and direct sums. A subset Φ of Spec A is specialization closed if for any pair 𝔭 ⊆ 𝔮 of prime ideals, 𝔭 ∈ Φ implies 𝔮 ∈ Φ.

Corollary 2.6 ([8, p. 425]). The map C ↦ Ass C induces a bijection between the set of localizing subcategories of Mod A and the set of specialization closed subsets of Spec A.

Proof. Suppose that C is localizing and let 𝔭 ⊆ 𝔮 be prime ideals with 𝔭 in Φ = Ass C. Then A/𝔭 ∈ C and therefore A/𝔮 ∈ C, because A/𝔮 is a factor module of A/𝔭. Thus 𝔮 ∈ Φ, and we have that Φ is specialization closed.

Now suppose that Φ ⊆ Spec A is specialization closed and let N ⊆ M be A-modules with M in C = Ass^{−1} Φ. Then M/N belongs to C, since

Ass M/N ⊆ {𝔭 ∈ Spec A | (M/N)_{𝔭} ≠ 0} ⊆ {𝔭 ∈ Spec A | M_{𝔭} ≠ 0} ⊆ Φ

where the last inclusion uses that Φ is specialization closed. Thus C is localizing. □

3. Subcategories via support

A full subcategory C of Mod A is called thick if for each exact sequence

M_{1} → M_{2} → M_{3} → M_{4} → M_{5}

of A-modules with M_{i} in C for i = 1, 2, 4, 5, the module M_{3} belongs to C. Note that a thick subcategory is an abelian category and that the inclusion functor is exact.

We wish to classify all thick subcategories of Mod A which are closed under taking direct sums. For this a few definitions are needed.

Let M be an A-module. Following [6], the support of M is by definition

Supp M = {𝔭 ∈ Spec A | Tor_{∗}^{A}(M, k(𝔭)) ≠ 0}.

For example, Supp I = Ass I for every injective A-module I.

Let Φ be a subset of Spec A and define the full subcategory

Inj_{Φ} A = {I ∈ Mod A | I is injective and Ass I ⊆ Φ}.

We call Φ coherent^{1} if each morphism I^{0} → I^{1} in Inj_{Φ} A can be completed to an exact sequence I^{0} → I^{1} → I^{2} with I^{2} in Inj_{Φ} A. For example, each specialization closed subset of Spec A is coherent.

Theorem 3.1. The map sending a subcategory C of Mod A to

Supp C = ⋃_{M∈C} Supp M

1The term coherent refers to the characterizing property of a coherent ring that every morphism P_{1} → P_{0} between finitely generated projective modules can be completed to an exact sequence P_{2} → P_{1} → P_{0} such that P_{2} is finitely generated projective.

induces a bijection between the set of full subcategories of Mod A, which are thick and closed under taking direct sums, and the set of coherent subsets of Spec A. The inverse map sends a subset Φ of Spec A to

Supp^{−1} Φ = {M ∈ Mod A | Supp M ⊆ Φ}.

Let us give an example of a set of prime ideals which is not coherent. This is based on an example from [2] and provides a counterexample to Hovey’s classification of thick subcategories closed under direct sums in [11, Theorem 5.2].

Example 3.2. Let k be a field and A = k[[x, y]]. Then Φ = Spec A\{𝔪} with 𝔪 = (x, y) is not coherent. To see this, let

0 → A → E(A) → ⨁_{ht𝔭=1}E(A/𝔭) → E(A/𝔪) → 0

be a minimal injective resolution of A. Then the morphism

E(A) → ⨁_{ht𝔭=1}E(A/𝔭)

cannot be completed to an exact sequence lying in Inj_{Φ} A, because its cokernel E(A/𝔪) does not belong to Inj_{Φ} A.

It should be clear that one can construct such examples more generally for commutative noetherian rings of Krull dimension at least two. On the other hand, if A is a Dedekind domain, then all subsets of Spec A are coherent, because every factor module of an injective A-module is injective and therefore the cokernel of a morphism I^{0} → I^{1} between injective A-modules is up to isomorphism a direct summand of I^{1}. We refer to Section 4 for details about coherent subsets.

The proof of Theorem 3.1 uses an alternative description of the support of a module. This involves the derived category D(Mod A) of Mod A. Given two complexes X and Y of A-modules, we write X ⊗_{A}^{L} Y for their tensor product in D(Mod A). Note that for every A-module M, we have

Supp M = {𝔭 ∈ Spec A | M ⊗_{A}^{L} k(𝔭) ≠ 0}.

The following lemma is due to Foxby. The proof given here is inspired by Neeman’s work [14, §2]; see Proposition 5.1 for a more general statement.

Lemma 3.3 ([6, Remark 2.9]). Let M be an A-module. Given a minimal injective resolution I^{∗} of M, we have

Supp M = ⋃_{i≥0} Ass I^{i}.

Proof. First observe that we can pass from M to the complex I, because the morphism M → I induces an isomorphism in D(Mod A). Fix a prime ideal 𝔭. Recall that each injective A-module J admits a unique decomposition

J = ⨁_{𝔮 prime}Γ_{𝔮}J

such that Ass Γ_{𝔮}J ⊆ {𝔮} for all 𝔮. We denote by Γ_{𝔭}I the complex which is obtained from I by taking in each degree the component with associated prime 𝔭. To be precise, Γ_{𝔭}I is the subcomplex of I_{𝔭} = I ⊗_{A} A_{𝔭} supported at the closed point 𝔭. Note that the sequence I → I_{𝔭} ← Γ_{𝔭}I of canonical morphisms is degreewise a split epimorphism, followed by a split monomorphism. In particular, it induces an isomorphism

I ⊗_{A}^{L} k(𝔭) ≅ I_{𝔭} ⊗_{A}^{L} k(𝔭) ≅ Γ_{𝔭}I ⊗_{A}^{L} k(𝔭),

since I′ ⊗_{A}^{L} k(𝔭) = 0 for the kernel I′ of I → I_{𝔭} and I′′ ⊗_{A}^{L} k(𝔭) = 0 for the cokernel I′′ of Γ_{𝔭}I → I_{𝔭}.

Suppose first that I ⊗_{A}^{L} k(𝔭) ≠ 0. Then Γ_{𝔭}I ≠ 0 and therefore 𝔭 ∈ Ass I^{i} for some i. Now suppose that I ⊗_{A}^{L} k(𝔭) = 0. Then Γ_{𝔭}I = 0 by [14, Lemma 2.14]. We want to conclude that Γ_{𝔭}(I^{i}) = (Γ_{𝔭}I)^{i} = 0 for all i. Here we need to use the minimality of I. Recall that a complex J of injective A-modules is minimal if for all i the kernel of the differential J^{i} → J^{i+1} is an essential submodule of J^{i}. If J is minimal and J^{i} = 0 for i ≪ 0, then H^{i}J = 0 for all i implies J^{i} = 0 for all i. Observe that Γ_{𝔭} preserves minimality. Thus Γ_{𝔭}I = 0 in D(Mod A) implies 𝔭 ∉ Ass I^{i} for all i, because I is minimal. □

Lemma 3.4. Let Φ be a coherent subset of Spec A. Then

Supp^{−1} Φ = {M ∈ Mod A | Supp M ⊆ Φ}

is a thick subcategory of Mod A.

Proof. We consider the full subcategory C consisting of all A-modules M which fit into an exact sequence

0 → M → I^{0} → I^{1} with Ass I^{i} ⊆ Φ (i = 0, 1).

Without any assumptions on Φ, it is clear that C is an additive subcategory of Mod A which is closed under taking kernels. An application of the horseshoe lemma shows that C is closed under forming extensions. Next observe that C is closed under taking cokernels. Here we use that Φ is coherent. By definition, the cokernel of a morphism between injective modules in C belongs to C. A standard argument then shows that this property extends to arbitrary morphisms in C. It follows from Lemma 3.3 that Supp^{−1} Φ = C, and therefore Supp^{−1} Φ is thick. □

Lemma 3.5. Let C be a subcategory of Mod A which is thick and closed under taking direct sums. Then the injective envelope E(M) belongs to C for every M in C.

Proof. Fix M in C. First observe that Tor_{i}^{A}(M, N) belongs to C for every A-module N and every integer i. This is clear, because for any projective resolution P of N, the complex M ⊗_{A} P and therefore its cohomology lies in C. Given 𝔭 ∈ Supp M, it follows that k(𝔭) belongs to C, since Tor_{i}^{R}(M, k(𝔭)) is a direct sum of copies of k(𝔭). Then Lemma 2.3 implies that E(A/𝔭) belongs to C, and we conclude from Lemma 3.3 that E(M) belongs to C. □

Proof of Theorem 3.1. Let Φ be a coherent subset of Spec A. Then the subcategory Supp^{−1} Φ is thick and closed under taking direct sums, by Lemma 3.4. Clearly, we have

Supp(Supp^{−1} Φ) = Φ.

Now let C be a subcategory of Mod A, which is thick and closed under taking direct sums. Let Φ = Supp C. First observe that Inj_{Φ} A ⊆ C, by Lemma 3.5. We claim that Φ is coherent. In deed, each morphism I^{0} → I^{1} in Inj_{Φ} A can be completed to an exact sequence I^{0} → I^{1} → I^{2} in Inj_{Φ} A by taking for I^{2} the injective envelope of a cokernel of I^{0} → I^{1}. Next we claim that

Supp^{−1}(Supp C) = C.

The inclusion Supp^{−1}(Supp C) ⊇ C is clear. Now suppose that M is a module contained in Supp^{−1}(Supp C) and choose a minimal injective resolution I^{∗}. Then Supp I^{i} ⊆ Supp C for all i. Thus I^{0} and I^{1} belong to C, and we conclude that M belongs to C. □

The classification of thick subcategories specializes to Gabriel’s classification of localizing subcategories.

Corollary 3.6 ([8, p. 425]). The map C ↦ Supp C induces a bijection between the set of localizing subcategories of Mod A and the set of specialization closed subsets of Spec A.

4. Coherent subsets of Spec A

In this section we collect some basic properties of coherent subsets of Spec A. Let us fix some notation. Given a multiplicatively closed subset S of A, let π: A → S^{−1}A denote the localization. Then we identify Spec S^{−1}A via π^{−1} with the subset of all prime ideals 𝔭 of A satisfying S ∩ 𝔭 = ∅.

Proposition 4.1. Let Φ be a subset of Spec A.

(1) Let (Φ_{i}) be a family of coherent subsets of Spec A. Then ⋂_{i} Φ_{i} is coherent.

(2) If Φ is specialization closed, then Φ is coherent.

(3) If 𝔮 ⊆ ⋃_{𝔭∈Φ} 𝔭 implies 𝔮 ∈ Φ for every prime ideal 𝔮, then Φ is coherent.

(4) The subset Φ is coherent if and only if Φ∩Spec A_{𝔭} is a coherent subset of Spec A_{𝔭} for each prime ideal 𝔭.

Proof. We use that Φ is coherent if and only if the cokernel C of each morphism I^{0} → I^{1} between injective A-modules with Ass I^{i} ⊆ Φ (i = 0, 1) satisfies Ass C ⊆ Φ.

(1) Clear.

(2) The assumption on Φ implies that for each pair N ⊆ M of A-modules with Ass M ⊆ Φ, we have that Ass M/N ⊆ Φ.

(3) The set

S = A \ ⋃_{𝔭∈Φ}𝔭 = ⋂_{𝔭∈Φ}A \ 𝔭

is multiplicatively closed. The assumption on Φ implies that the localization A → S^{−1}A identifies all injective S^{−1}A-modules with the injective A-modules I satisfying Ass I ⊆ Φ.

(4) We write Φ_{𝔭} = Φ∩Spec A_{𝔭} for each prime ideal 𝔭. Suppose first that Φ is coherent and fix a prime ideal 𝔭. Let I^{0} → I^{1} be a map in Inj_{Φ_{𝔭}} A. There exist an exact sequence I^{0} → I^{1} → I^{2} in Inj_{Φ} A and localization at 𝔭 induces an exact sequence in Inj_{Φ_{𝔭}} A. Thus Φ_{𝔭} is coherent. Now suppose that Φ is not coherent. It follows that there exists an exact sequence

I^{0} → I^{1} → C → 0

of A-modules with I^{i} ∈ Inj_{Φ} A (i = 0, 1) but Ass C ⊈ Φ. Let 𝔭 ∈ Ass C \ Φ. Then we localize at 𝔭 and obtain an exact sequence

I_{𝔭}^{0} → I_{𝔭}^{1} → C_{𝔭} → 0

of A_{𝔭}-modules with I_{𝔭}^{i} ∈ Inj_{Φ_{𝔭}} A (i = 0, 1) but

Ass C_{𝔭} = (Ass C) ∩ Spec A_{𝔭} ⊈ Φ_{𝔭}.

Thus Φ_{𝔭} is a subset of Spec A_{𝔭} which is not coherent. □

Remark 4.2. (1) A subset Φ of Spec A satisfies the condition (3) of Proposition 4.1 if and only if it is of the form Spec S^{−1}A for some multiplicatively closed subset S.

(2) The union of two coherent subsets need not to be coherent. For instance, Example 3.2 provides a subset which is not coherent but Zariski open. Each Zariski open subset U can be written as the finite union of basic open subsets. However, a basic open set is coherent because it is of the form Spec S^{−1}A for some multiplicatively closed subset S.

Corollary 4.3. If the Krull dimension of A is at most one, then every subset of Spec A is coherent.

The converse of this statement is proved in the appendix of this paper.

Proof. We may assume that A is local, by part (4) of Proposition 4.1, and we denote by 𝔪 the maximal ideal. Let Φ be a subset of Spec A. If Φ contains 𝔪, then Φ is specialization closed and therefore coherent, by part (2) of Proposition 4.1. If 𝔪 is not contained in Φ, then all prime ideals in Φ are minimal and therefore the prime avoidance theorem implies that the condition in part (3) of Proposition 4.1 is satisfied. Thus Φ is coherent. □

Given a prime ideal 𝔭 of A, let

V (𝔭) = {𝔮 ∈ Spec A | 𝔭 ⊆ 𝔮} and Λ(𝔭) = {𝔮 ∈ Spec A | 𝔮 ⊆ 𝔭}.

Subsets of the from V (𝔭) and Λ(𝔭) are coherent. They can be used to build new coherent subsets.

Corollary 4.4. Let Φ and Ψ be subsets of Spec A and suppose that Ψ is finite. Then

⋃_{𝔭∈Φ,𝔮∈Ψ}V (𝔭) ∩ Λ(𝔮)

is a coherent subset of Spec A.

Proof. We can express the set as the intersection of two coherent subsets:

⋃_{𝔭∈Φ,𝔮∈Ψ}V (𝔭) ∩ Λ(𝔮) = ( ⋃_{𝔭∈Φ}V (𝔭)) ∩ ( ⋃_{𝔮∈Ψ}Λ(𝔮)).  □

Example 4.5. Let 𝔭_{1}, 𝔭_{2} be prime ideals. If 𝔭_{1} ⊆ 𝔭_{2}, then {𝔮 | 𝔭_{1} ⊆ 𝔮 ⊆ 𝔭_{2}} is coherent. If 𝔭_{1} ⊈ 𝔭_{2} and 𝔭_{2} ⊈ 𝔭_{1}, then {𝔭_{1}, 𝔭_{2}} is coherent. In both cases the set is of the form

⋃_{𝔭,𝔮∈Φ}V (𝔭) ∩ Λ(𝔮) with Φ = {𝔭_{1}, 𝔭_{2}}.

5. Support of complexes

Let X be a complex of A-modules. Following [6], the support of X is by definition

Supp X = {𝔭 ∈ Spec A | X ⊗_{A}^{L} k(𝔭) ≠ 0}.

We use an alternative description of the support of X. A complex I of injective Amodules together with a quasi-isomorphism X → I is called a minimal K-injective resolution of X, if I is K-injective (that is, every morphism from an acyclic complex to I is null homotopic) and I is minimal (that is, for all i the kernel of the differential I^{i} → I^{i+1} is an essential submodule of I^{i}).

One can show that each complex of A-modules admits a minimal K-injective resolution; see [15, Theorem 4.5] or [3, Application 2.4] for the existence of a K-injective resolution and [12, Proposition B.2] for the minimality. Note that each acyclic and K-injective complex is null homotopic. Moreover, if a minimal complex I of injective A-modules is null homotopic, then I^{i} = 0 for all i.

The next proposition is the obvious generalization of Lemma 3.3 from modules to complexes of modules. The proof requires only minor modifications; it follows closely [14, §2].

Proposition 5.1. Let X be a complex of A-modules and X → I a minimal K-injective resolution of X. Then we have

Supp X = ⋃_{i∈ℤ}Ass I^{i}.

Proof. First observe that we can pass from X to the complex I, because the morphism X → I induces an isomorphism in D(Mod A). Fix a prime ideal 𝔭. Recall that each injective A-module J admits a unique decomposition

J = ⨁_{𝔮 prime}Γ_{𝔮}J

such that Ass Γ_{𝔮}J ⊆ {𝔮} for all 𝔮. We denote by Γ_{𝔭}I the complex which is obtained from I by taking in each degree the component with associated prime 𝔭. To be precise, Γ_{𝔭}I is the subcomplex of I_{𝔭} = I ⊗_{A} A_{𝔭} supported at the closed point 𝔭. Note that the sequence I → I_{𝔭} ← Γ_{𝔭}I of canonical morphisms is degreewise a split epimorphism, followed by a split monomorphism. In particular, it induces an isomorphism

I ⊗_{A}^{L} k(𝔭) ≅ I_{𝔭} ⊗_{A}^{L} k(𝔭) ≅ Γ_{𝔭}I ⊗_{A}^{L} k(𝔭),

since I′ ⊗_{A}^{L} k(𝔭) = 0 for the kernel I′ of I → I_{𝔭} and I′′ ⊗_{A}^{L} k(𝔭) = 0 for the cokernel I′′ of Γ_{𝔭}I → I_{𝔭}.

Suppose first that I ⊗_{A}^{L} k(𝔭) ≠ 0. Then Γ_{𝔭}I ≠ 0 and therefore 𝔭 ∈ Ass I^{i} for some i. Now suppose that I ⊗_{A}^{L} k(𝔭) = 0. Then Γ_{𝔭}I = 0 by [14, Lemma 2.14], that is, Γ_{𝔭}I is acyclic. We want to conclude that Γ_{𝔭}(I^{i}) = (Γ_{𝔭}I)^{i} = 0 for all i. Here we need to use the minimality of I. We observe that Γ_{𝔭} preserves minimality. Also, Γ_{𝔭}I is a K-injective complex of injective A-modules. Thus Γ_{𝔭}I = 0 in D(Mod A) implies 𝔭 ∉ Ass I^{i} for all i, because I is minimal. □

Theorem 5.2. For a subset Φ of Spec A the following conditions are equivalent: (1) Φ is coherent.

(2) For every complex X of A-modules we have

Supp X ⊆ Φ ⟺ Supp H^{i}X ⊆ Φ for all i ∈ ℤ.

(3) For every complex X of A-modules we have

Supp X ⊆ Φ ⟹ Supp H^{i}X ⊆ Φ for all i ∈ ℤ.

Proof. (1) ⇒ (2): Suppose that Φ is coherent. We use that the A-modules M with Supp M ⊆ Φ form a thick subcategory, by Lemma 3.4. Now fix a complex X of Amodules. If Supp X ⊆ Φ, then we have a quasi-isomorphic complex I of injective A-modules with Ass I^{i} ⊆ Φ for all i, by Proposition 5.1. It follows that

Supp H^{i}X = Supp H^{i}I ⊆ Φ

for all i. Now let Supp H^{∗}X ⊆ Φ and 𝔭 ∈ Supp X. A dévissage argument shows that Supp H^{∗}(X ⊗_{A}^{L} Y) ⊆ Φ for every complex Y. In particular,

{𝔭} = Supp H^{∗}(X ⊗_{A}^{L} k(𝔭)) ⊆ Φ

because X ⊗_{A}^{L} k(𝔭) is a direct sum of shifted copies of k(𝔭). Thus Supp X ⊆ Φ. (2) ⇒ (3): Clear.

(3) ⇒ (1): Let I^{0} → I^{1} be a morphism of injective A-modules with Ass I^{i} ⊆ Φ for i = 0, 1. Viewing I as a complex, we have Supp I ⊆ Φ by Proposition 5.1, and therefore Supp H^{0}I ⊆ Φ. It follows from Lemma 3.3 that we can complete I^{0} → I^{1} to an injective resolution

0 → H^{0}I → I^{0} → I^{1} → I^{2} → ⋯

of H^{0}I with Ass I^{i} ⊆ Φ for all i. Thus Φ is coherent. □

Acknowledgements. A pleasant collaboration with Dave Benson and Srikanth Iyengar has been the starting point of this paper; in particular the Example 3.2 is taken from [2]. I would like to thank Amnon Neeman and Torsten Wedhorn for a number of helpful comments on this work.

Appendix A. Noncoherent subsets of Spec A

By Srikanth Iyengar

In this appendix, we establish the converse of Corollary 4.3. To this end, we recall some standard notions from commutative algebra; this serves also to fix notation.

Definition A.1. Let A be a commutative noetherian ring, 𝔞 an ideal in A, and let M be an A-module. The 𝔞-depth of M is the number

depth_{A}(𝔞, M) = inf{n | Ext_{A}^{n}(A/𝔞, M) ≠ 0} .

This invariant of M can also be detected from its Koszul homology on a finite set of elements generating 𝔞, and also its local cohomology with respect to 𝔞; see, for instance, [7, Theorem 2.1]. When M is finitely generated and 𝔞M ≠ M, this number coincides with the length of the longest M-regular sequence in 𝔞; see [13, Theorem 28].

As usual, if A is local, with maximal ideal 𝔪, we write depth_{A} M for the 𝔪-depth of M, and call it the depth of M.

We record the following standard properties of depth, for ease of reference.

Lemma A.2. Let 𝔞 be an ideal in a commutative noetherian ring A, and let M be an A-module. The following statements hold.

(1) One has an equality, depth_{A}(\sqrt{𝔞}, M) = depth_{A}(𝔞, M).

(2) With I^{∗} the minimal injective resolution of M, for each prime ideal 𝔭 one has

depth_{A_{𝔭}} M_{𝔭} = inf{n | E(A/𝔭) is a direct summand in I^{n}} .

(3) If A → B is a homomorphism of rings and N is a B-module, then viewing N as an A-module by restriction of scalars, one has

depth_{A}(𝔞, N) = depth_{B}(𝔞B, N) .

Proof. For (1) see, for instance, [7, Proposition 2.11], while (3) is evident, if one computes depth using Koszul complexes, or via local cohomology; see [7, Theorem 2.1]. Part (2) holds as (I^{∗})_{𝔭} is the minimal injective resolution of M_{𝔭} over A_{𝔭}, so the complex Hom_{A_{𝔭}}(A_{𝔭}/𝔭A_{𝔭}, (I^{∗})_{𝔭}), whose first nonzero cohomology module occurs in degree depth_{A_{𝔭}} M_{𝔭}, has zero differential. □

We need the following result of Auslander and Buchsbaum, which is implicit in [1].

Proposition A.3. Let A be a commutative noetherian ring with dim A ≥ 1. There exists a prime 𝔭 in Spec A such that

depth A_{𝔭} = dim A_{𝔭} = dim A − 1.

Proof. The proof uses an induction on dim A. When dim A = 1, for 𝔭 one may take any minimal prime of A. This is the basis of the induction.

Assume dim A ≥ 2, and let 𝔪 be the maximal ideal of A. Since Ass A is finite, the prime avoidance theorem implies that the following set in nonempty:

𝔪 \ ⋃_{𝔭≠𝔪^{𝔭∈Ass A}} 𝔭 .

Choose an element x in it. One then has that dim(A/Ax) = dim A − 1, so the induction hypothesis yields a prime 𝔭 in A, containing x, such that

depth(A_{𝔭}/A_{𝔭}x) = dim(A_{𝔭}/A_{𝔭}x) = dim(A/Ax) − 1 = dim A − 2 .

Observe that 𝔭 ≠ 𝔪, so the choice of x ensures that it is a nonzero divisor in A_{𝔭}. Thus one has depth A_{𝔭} = dim A_{𝔭} = dim A − 1. This completes the induction argument. □

The gist of the result below is well-known; we provide a proof for lack of a suitable reference for this formulation.

Theorem A.4. Let A be a commutative noetherian ring with dim A ≥ 2. There exists an A-module M and a prime 𝔭 in Spec A such that depth_{A_{𝔭}} M_{𝔭} = max{2, dim A − 1}.

Proof. When dim A ≥ 3, we apply Proposition A.3, and take for 𝔭 any prime such that depth A_{𝔭} = dim A − 1 and set M = A. Henceforth, we assume dim A = 2. Choosing a prime ideal 𝔭 in A with dim A_{𝔭} = 2, and replacing A with A_{𝔭}, we may assume also that A is local; the goal then is to find an A-module M such that depth_{A} M = 2. At this point, one may refer to, for instance, Hochster’s article [10], especially Section 3. We provide details, for completeness.

Let Â be the completion of A at its maximal ideal, say 𝔪. One has dim Â = 2, so there exists a prime ideal 𝔞 in Â with dim(Â/𝔞) = 2. Consider the canonical homomorphisms of rings A → Â → Â/𝔞. Observer that 𝔪(Â/𝔞) is the maximal ideal of Â/𝔞, so, by Lemma A.2(3), for any module M over Â/𝔞, one has

depth_{A} M = depth_{Â/𝔞} M .

Replacing A with Â/𝔞 one may assume A is a complete local domain, with dim A = 2.

Let B be the integral closure of A in its field of fractions; the conditions on A imply that B is finite as an A-module, by [13, Corollary 2, pp. 234], so also a two dimensional noetherian ring, by [13, Theorem 20]. The finiteness of the extension A ⊆ B implies that there exists a prime ideal 𝔮 in B with dim B_{𝔮} = 2 and 𝔮 ∩ A = 𝔪, the maximal ideal of A; see [13, Theorem 5(iii)]. The choice of 𝔮 ensures that \sqrt{𝔪B_{𝔮}} = 𝔮B_{𝔭}, so part (3) and (2) of Lemma A.2 yield the first and second equalities below:

depth_{A} B_{𝔮} = depth_{B_{𝔮}}(𝔪B_{𝔮}, B_{𝔮})

= depth B_{𝔮} = dim B_{𝔮} = 2

The penultimate equality holds because B_{𝔮} is Cohen-Macaulay, by Serre’s criterion for normality; see [13, Theorem 39]. The A-module B_{𝔮} has thus the desired depth. □

The result below is a perfect converse to Corollary 4.3.

Corollary A.5. If A is a commutative noetherian ring with dim A ≥ 2, then there exists a subset Φ of Spec A that is not coherent.

Proof. By Theorem A.4, there exists an A-module M and a prime ideal 𝔭 in A such that d = depth_{A_{𝔭}} M_{𝔭} ≥ 2. Let I^{∗} be a minimal injective resolution of M, and set

Φ = Ass_{A}(I^{d−2}) ∪ Ass_{A}(I^{d−1})

Observe that 𝔭 is in Ass_{A}(I^{d}) but not in Φ, by Lemma A.2(2). Thus, the set Φ is not coherent, as the associated primes of Coker(I^{d−2} → I^{d−1}) coincide with those of I^{d}. □

References

[1] M. Auslander and D. Buchsbaum: Homological dimension in noetherian rings. II. Trans. Amer. Math. Soc. 88 (1958), 194–206.

[2] D. Benson, S. Iyengar, and H. Krause: Local cohomology and support for triangulated categories. arXiv:math.KT/0702610.

[3] M. Bökstedt and A. Neeman: Homotopy limits in triangulated categories. Compositio Math. 86 (1993) 209–234.

[4] N. Bourbaki: Algèbre commutative. Hermann, Paris, 1968.

[5] N. Bourbaki: Algèbre commutative. Chapitre 10. Masson, Paris, 1998.

[6] H.-B. Foxby: Bounded complexes of flat modules. J. Pure Appl. Algebra 15 (1979), 149–172.

[7] H.-B. Foxby and S. Iyengar: Depth and amplitude for unbounded complexes. Commutative algebra (Grenoble/Lyon, 2001), 119–137, Contemp. Math., 331, Amer. Math. Soc., Providence, RI, 2003.

[8] P. Gabriel: Des catégories abéliennes. Bull. Soc. Math. France 90 (1962), 323–448.

[9] G. Garkusha and M. Prest: Classifying Serre subcategories of finitely presented modules. Preprint (2006).

[10] M. Hochster: Cohen-Macaulay modules. Conference on Commutative Algebra (Lawrence, Kansas 1972), 120–152, Lecture Notes Math. 311, Springer-Verlag, 1973.

[11] M. Hovey: Classifying subcategories of modules. Trans. Amer. Math. Soc. 353 (2001), 3181–3191.

[12] H. Krause: The stable derived category of a noetherian scheme. Compos. Math. 141 (2005), 1128–1162.

[13] H. Matsumura: Commutative algebra. Second edition. Mathematics Lecture Note Series, 56, Benjamin/Cummings Publishing Co., Inc., Reading, Mass., 1980.

[14] A. Neeman: The chromatic tower of D(R). Topology 31 (1992), 519–532.

[15] N. Spaltenstein: Resolutions of unbounded complexes. Compositio Math. 65 (1988), 121–154.

[16] R. Takahashi: Classifying subcategories of modules over commutative noetherian rings. Preprint (2006).

Srikanth Iyengar, Department of Mathematics, University of Nebraska, Lincoln NE 68588, U.S.A.

   E-mail address: iyengar@math.unl.edu

Henning Krause, Institut für Mathematik, Universität Paderborn, 33095 Paderborn, Germany.

   E-mail address: hkrause@math.upb.de